OpenStudy (anonymous):

Seven slips of paper marked with the numbers 1, 2, 3, 4, 5, 6, and 7 are placed in a box and mixed well. Two are drawn. What are the odds that the sum of the numbers on the two slips is 8?

OpenStudy (anonymous):

=)

OpenStudy (anonymous):

You probably better just list them out.

OpenStudy (anonymous):

There are 1+7 2+6 3+5 ---- 3 ways to make 8 and there are 7 ways to draw the first number, and 6 ways to draw the second ways to make 8 ------------- = 3/(7*6)=1/14 total ways to draw If the order of drawing them matter then there are twice as many ways to make 8 that is, if 1+7 is distinct from 7+1, that would double the number of ways to make 8, which is an answer choice.

OpenStudy (anonymous):

Order is distinct in this case.

OpenStudy (anonymous):

Doesnt specify if order is important. One of the answer choices shows 1 to 7 but the test doesnt tell me if its correct or not. :(

OpenStudy (anonymous):

use your brain... Clearly if someone drew 1, 7 it is different than drawing 7, 1

OpenStudy (anonymous):

I choose 1 to 7. All i was saying is that the test doenst tell me if it was correct or not. Thanks